OpenStudy (anonymous):

Solve the equation on the interval 0

OpenStudy (campbell_st):

well find the inverse value \[2x = \sin^{-1}(\frac{1}{2})\] then solve for x and given the domain, you'll need to identify angles in the 1st and 2nd quadrants

OpenStudy (anonymous):

sin 2x = 1/2 The trig conversion table gives sin Pi/6 = 1/2 and sin 5Pi/6 = 1/2 Two answers: a) sin 2x = sin Pi/6 -> 2x = Pi/6 -> x = Pi/12 b) sin 2x = sin 5Pi/6 -> 2x = 5Pi/6 --> x = 5Pi/12